# Conditioning of Banach Space Valued Gaussian Random Variables: An Approximation Approach Based on Martingales

### v1

@@ B Appendix: Covariance Operators
In other words, the cross covariance operator provides the covariances between all pairs of one-dimensional projections of $X$ and $Y$. Another routine check shows that the dual is given by
@@
Indeed, (B.3) follows by applying (A.10) with the operator $S$ and using the definition of $T^{\prime}$ and (B.4) is a direct consequence of the definition of $\operatorname{cov}(X,Y)$.
@@
for all $x^{\prime},x_{1}^{\prime},x_{2}^{\prime}\in E^{\prime}$, where $\operatorname{var}(Z)$ denotes the variance of an $\mathbb{R}$-valued random variable $Z$. This motivates the following definition.
@@
for all $x_{1}^{\prime},x_{2}^{\prime}\in E^{\prime}$. The following lemma uses this observation to show that abstract covariance operators are determined by the behavior of $x^{\prime}\mapsto\langle x^{\prime},Tx^{\prime}\rangle_{E^{\prime},E}$.
@@
where the second identity can be found in e.g. [48, Theorem 1.1.2]. Moreover, a straightforward calculation shows that $\Lambda^{\dagger}$ is the diagonal $n\times n$-matrix having the entries $\lambda_{1}^{-1},\dots,\lambda_{n}^{-1}$ with $00^{-1}:=$ 0. Consequently, $A^{\dagger}$ is also symmetric and positive semi-definite. Moreover, this matrix describes the operator $T^{\dagger}:\mathbb{R}^{n}\to(\mathbb{R}^{n})^{\prime}$, and therefore $T^{\dagger}$ is symmetric and non-negative.
@@
Lemma B.2 shows that this kernel is determined by its behaviour on the diagonal. Moreover, the Cauchy-Schwarz inequality for kernels, see e.g. [38, (4.14)] yields
@@
In particular, $\langle x_{2}^{\prime},Tx_{2}^{\prime}\rangle_{E^{\prime},E}=0$ implies $Tx_{2}^{\prime}=0$ by a simple application of Hahn-Banach’s theorem.
@@
Let us therefore fix some $x^{\prime}_{0}\in B_{E^{\prime}}$ as well as a sequence $(x_{n}^{\prime})\subset B_{E^{\prime}}$ such that $x_{n}^{\prime}\to x^{\prime}_{0}$ with respect to $w^{*}$. Moreover, we write $Z:=X-\mathbb{E}X$. By (B.4) and (B.1) we then find
@@
Let $E$ and $F$ be separable Banach spaces, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $X:\Omega\to E$ and $Y:\Omega\to F$ be centered Gaussian random variables, and $S:F\to F^{\prime}$ be a bounded, linear, symmetric, and non-negative operator. Then the operator $K:E^{\prime}\to E$ defined by
@@
$\langle x_{2}^{\prime},Kx_{1}^{\prime}\rangle_{E^{\prime},E}$ $=\langle x_{2}^{\prime},\operatorname{cov}(X,Y)y_{1}^{\prime}\rangle_{E^{\prime},E}$
@@
$=\int_{\Omega}\int_{\Omega}\langle x_{1}^{\prime},X(\omega_{1})\rangle_{E^{\prime},E}\cdot\langle S\circ Y(\omega_{1}),Y(\omega_{2})\rangle_{F^{\prime},F}\cdot\langle x_{2}^{\prime},X(\omega_{2})\rangle_{E^{\prime},E}\,\mathrm{d}\mathrm{P}(\omega_{1})\,\mathrm{d}\mathrm{P}(\omega_{2})\,.$
@@
for all $\omega_{1},\omega_{2}\in\Omega$. Now Fernique’s theorem ensures $\|X\|_{E},\|Y\|_{F}\in{{\cal L}_{2}(\mathrm{P})}$, and therefore we find $\|X\|_{E}\cdot\|Y\|_{F}\in{{\cal L}_{1}(\mathrm{P})}$ by the Cauchy-Schwarz inequality. This gives
@@
$\leq\|x_{1}^{\prime}\|\cdot\|x_{2}^{\prime}\|\cdot\|S\|\int_{\Omega}\|X(\omega_{1})\|_{E}\cdot\|Y(\omega_{1})\|_{F}\,\mathrm{d}\mathrm{P}(\omega_{1})\cdot\int_{\Omega}\|X(\omega_{2})\|_{E}\cdot\|Y(\omega_{2})\|_{F}\,\mathrm{d}\mathrm{P}(\omega_{2})$
@@
and therefore Fubini’s theorem yields the formula for $\langle x_{2}^{\prime},Kx_{1}^{\prime}\rangle_{E^{\prime},E}$.
@@
$(x^{\prime}\otimes y)(x):=\langle x^{\prime},x\rangle_{E^{\prime},E}\cdot y\,,\qquad\qquad x\in E\,.$
@@
for all $x^{\prime},x^{\prime}_{1},x^{\prime}_{2}\in E$, $y,y_{1},y_{2}\in F$, and $\alpha,\beta\in\mathbb{R}$, where, as usual, we assumed that $\otimes$ has a higher priority than $+$. We define the algebraic tensor product space by
@@
It is well-known that the set $\mathfrak{N}(E,F)$ of nuclear operators $T:E\to F$ equipped with $\|\,\cdot\,\|_{\mathrm{nuc}}$ is a Banach space and we have $\|T\|\leq\|T\|_{\mathrm{nuc}}$ for all $T\in\mathfrak{N}(E,F)$, see e.g. [50, Satz VI.5.3] or [12, Theorem 5.25]. Moreover, every finite rank operator is nuclear, and in particular all elementary tensors are nuclear. In addition, for $x^{\prime}\in E^{\prime}$ and $y\in F$ we have
@@
Under suitable assumptions, it turns out that the Bochner integral of tensor space valued functions can be used to compute cross covariances. To establish these results we need the following elementary lemma, which investigates such functions in the case of $X$ or $Y$ being constant.
@@
Proof of Lemma B.6: We only prove the assertion for $X^{\prime}\otimes y$, since the proof for $x^{\prime}\otimes Y$ is analogous. Let us first assume that $X^{\prime}$ is a measurable step function. We fix a representation $X=\sum_{i=1}^{m}\boldsymbol{1}_{A_{i}}x_{i}^{\prime}$, where $A_{i}\in{\cal A}$ and $x_{i}^{\prime}\in E^{\prime}$. For $\omega\in\Omega$ the bilinearity of elementary tensors then gives
@@
$\|(X_{n}^{\prime}\otimes y)(\omega)-(X^{\prime}\otimes y)(\omega)\|_{\mathrm{nuc}}\to 0\,.$
Furthermore, (A.2) and (B.17) gives $\|(X^{\prime}_{n}\otimes y)(\omega)\|_{\mathrm{nuc}}\leq\|(X^{\prime}\otimes y)(\omega)\|_{\mathrm{nuc}}$, and we already know that each $X^{\prime}_{n}\otimes y:\Omega\to\mathfrak{N}(E,F)$ is a measurable step function. Consequently, $X^{\prime}\otimes y:\Omega\to\mathfrak{N}(E,F)$ is strongly measurable as discussed around (A.1) and (A.2). Moreover, with the help of (B.16) we obtain
@@
Let us now turn to the announced integral representation of cross covariance operators. To this end, we fix two Banach spaces $E$ and $F$, as well as $x\in E$ and $y\in F$. For $x^{\prime\prime}:=\iota_{E}x\in E^{\prime\prime}$, where $\iota_{E}:E\to E^{\prime\prime}$ denotes the canonical, isometric embedding of $E$ into its bidual $E$, we then write
@@
$\|x\otimes y\|$ $=\|x\otimes y\|_{\mathrm{nuc}}=\|x\|_{E}\cdot\|y\|_{F}\,,$
$\|x_{1}\otimes y_{1}\|_{\mathrm{nuc}}$ $\leq\|x_{2}\otimes y_{2}\|_{\mathrm{nuc}}$
for all $x,x_{1},x_{2}\in E$ and $y,y_{1},y_{2}\in F$ with $\|x_{1}\|_{E}\leq\|x_{2}\|_{E}$ and $\|y_{1}\|_{F}\leq\|y_{2}\|_{F}$. Moreover, Lemma B.5 also holds, since $x_{n}\to x$ in $E$ implies $\iota_{E}x_{n}\to\iota_{E}x$ in $E^{\prime\prime}$.
@@
$\int_{\Omega}X\otimes y\,\mathrm{d}\mathrm{P}=\int_{\Omega}X^{\prime\prime}\otimes y\,\mathrm{d}\mathrm{P}=\biggr{(}\int_{\Omega}X^{\prime\prime}\,\mathrm{d}\mathrm{P}\biggl{)}\otimes y=\biggr{(}\int_{\Omega}\iota_{E}\circ X\,\mathrm{d}\mathrm{P}\biggl{)}\otimes y$ $=\biggl{(}\iota_{E}\biggr{(}\int_{\Omega}X\,\mathrm{d}\mathrm{P}\biggl{)}\biggr{)}\otimes y$
$=\biggl{(}\int_{\Omega}X\,\mathrm{d}\mathrm{P}\biggr{)}\otimes y$
by an application of (A.10) and our nomenclature for tensors of the form $x\otimes y$. In particular, the Bochner integral of $X$ is an $E$-valued Bochner integral and not just an $E^{\prime\prime}$-valued integral as a naïve application of Lemma B.6 would suggest.
@@
where in the second to last step we used Lemma B.6 twice.
@@
Furthermore, (A.2) and (B.17) gives $\|(X_{n}\otimes Y_{n})(\omega)\|_{\mathrm{nuc}}\leq\|(X\otimes Y)(\omega)\|_{\mathrm{nuc}}$. In addition, we already know that each $X_{n}\otimes Y_{n}:\Omega\to\mathfrak{N}(E^{\prime},F)$ is a measurable step function. Consequently, $X\otimes Y$ is strongly measurable as discussed around (A.1) and (A.2). Moreover, with the help of (B.16) we obtain
@@
To establish (B.18), we first note that (A.1) and (A.2) for $X$ and $(X_{n})$ give $\mathbb{E}X_{n}\to\mathbb{E}X=0$ by the definition of Bochner integrals and our assumption that $X$ is centered. Analogously, we find $\mathbb{E}Y_{n}\to 0$. Combining both gives $\mathbb{E}X_{n}\otimes\mathbb{E}Y_{n}\to 0$ by Lemma B.5. For $x^{\prime}\in E^{\prime}$ we thus find
@@
where in the first step we used the definition of Bochner integrals, the second step used that the limit in the second expression is with respect to the nuclear norm and thus also with respect to the operator norm, see also the explanation following (A.4), and in the last step we used the already established identity (B.18) for step functions. Moreover, we also have
@@
where the second step is based upon the dominated convergence property discussed around (A.7) to (A.9) with
@@
for all $\omega\in\Omega$ and the $\mathrm{P}$-integrable $g:\Omega\to[0,\infty]$ defined by $g(\omega):=\|x^{\prime}\|_{E^{\prime}}\cdot\|X(\omega)\|_{E}\cdot\|Y(\omega)\|_{F}$. Combining (B.19) and (B.20) yields (B.18). ∎
@@ C Appendix: Characteristic Functions
It is well known, see e.g. [27, Theorem 15.21], that every characteristic function $\varphi_{X}$ is uniformly continuous and we have $|\varphi_{X}(t)|\leq 1$ for all $t\in\mathbb{R}$ with $\varphi_{X}(0)=1$. Moreover, the characteristic function can be used to identify measures. The following well-known theorem, which can be found in e.g. [10, Proposition 10.3.11] or [27, Theorem 15.8] formalizes this.
@@
Proof of Lemma C.2: The second identity can be found in e.g. [22, Theorem 13.3], and the first identity is a special case of the second since $\langle a,X\rangle=a^{\mathsf{T}}X$ and $as=sa$, where $as$ denotes the matrix product and $a$ are $s$ are viewed as $d\times 1$ and $1\times 1$-matrices, respectively. ∎
@@
iii) $\Rightarrow$ i). See, for example, [14, Theorem 9.8.2] or [27, Theorem 15.23].
ii) $\Rightarrow$ iii). Let us fix a $t\in\mathbb{R}^{d}$ with $t\neq 0$. We write $a:=\|t\|_{2}^{-1}t$ and $s:=\|t\|_{2}$ as well as $Y_{n}:=\langle a,X_{n}\rangle$ and $Y:=\langle a,X\rangle$. By Lemma C.2 we then obtain
$\varphi_{X_{n}}(t)=\varphi_{X_{n}}(sa)=\varphi_{\langle a,X_{n}\rangle}(s)=\varphi_{Y_{n}}(s)$
and analogously we find $\varphi_{X}(t)=\varphi_{Y}(s)$. Moreover, we have $Y_{n}\to Y$ in distribution and hence the already established implication i) $\Rightarrow$ iii) gives $\varphi_{Y_{n}}(s)\to\varphi_{Y}(s)$. The previous two identities then yield $\varphi_{X_{n}}(t)\to\varphi_{X}(t)$ for all $t\neq 0$. Moreover, we always have $\varphi_{X_{n}}(0)=1=\varphi_{X}(0)$, and hence iii) has been shown.
i) $\Rightarrow$ ii). Let us fix a bounded and continuous $f:\mathbb{R}\to\mathbb{R}$. Then $f(\langle a,\,\cdot\,\rangle):\mathbb{R}^{d}\to\mathbb{R}$ is also bounded and continuous, and hence the transformation formula for image measures together with i) yields
$\int_{\mathbb{R}}f\,\mathrm{d}\mathrm{P}_{\langle a,X_{n}\rangle}=\int_{\mathbb{R}^{d}}f(\langle a,\,\cdot\,\rangle)\,\mathrm{d}\mathrm{P}_{X_{n}}\to\int_{\mathbb{R}^{d}}f(\langle a,\,\cdot\,\rangle)\,\mathrm{d}\mathrm{P}_{X}=\int_{\mathbb{R}}f\,\mathrm{d}\mathrm{P}_{\langle a,X\rangle}\,,$
that is, we have found ii). ∎
@@
Since $\varphi_{X}$ is uniformly continuous with $\varphi_{X}(0)=1$, there exists a $t_{0}>0$ such that $a_{t}>0$ for all $t\in[-t_{0},t_{0}]$. Consequently, we have
@@
see e.g. [46, Chapter IV.2.1]. Like in the finite dimensional case, the characteristic function can be used to identify probability measures. Namely, if $\mathrm{P}$ and $\mathrm{Q}$ are probability measures with $\varphi_{\mathrm{P}}=\varphi_{\mathrm{Q}}$, then we have $\mathrm{P}=\mathrm{Q}$, see e.g. [46, Theorem 2.2 in Chapter IV.2.1].
Similar to the finite dimensional case, a Gaussian measure $\mathrm{P}$ on a separable Banach space $E$ has a characteristic function that is solely described by its mean and covariance. The following result, which can be found in e.g. [46, Proposition 2.8 in Chapter IV.2.4], provides the details.
@@
Proof of Theorem C.6: We will apply [46, Theorem 3.1 in Chapter IV.3.1] for $\Gamma:=E^{\prime}$. Here we first note that Hahn-Banach’s theorem ensures that $E^{\prime}$ is a separating set consisting of continuous functions $E\to\mathbb{R}$. Moreover, every probability measure on $\mathfrak{B}(E)$ is a Radon probability measure, see the references around (44). Finally, [6, Theorem 3.8.4] in combination with i) ensures that $\{\mathrm{P}_{n}:n\geq 1\}$ is relatively weakly compact since the underlying space $E$ is a metric space, and thus a completely regular topological space, where the latter notion can be found in e.g. [25, Chapter 4 on p. 117], see also [6, Appendix A.1 on page 363]. Now the assertion follows from [46, Theorem 3.1 in Chapter IV.3.1]. ∎

### v3

@@ B Appendix: Covariance Operators
In other words, the cross covariance operator provides the covariances between all pairs of one-dimensional projections of $X$ and $Y$. In particular, if $X$ and $Y$ are independent, then all pairs $\langle x^{\prime},X\rangle_{E^{\prime},E}$ and $\langle y^{\prime},Y\rangle_{F^{\prime},F}$ are independent, and hence we find $\operatorname{cov}(X,Y)=0$. Another routine check shows that the dual $\operatorname{cov}(X,Y)^{\prime}:E^{\prime}\to F^{\prime\prime}$ of $\operatorname{cov}(X,Y)$ is given by
@@
Indeed, (B.3) follows by applying (A.10) with the operator $S$ and using the definition of $T^{\prime}$. Moreover, (B.4) is a direct consequence of the definition of $\operatorname{cov}(X,Y)$.
@@
for all $x^{\prime},x_{1}^{\prime},x_{2}^{\prime}\in E^{\prime}$, where $\operatorname{var}(Z)$ denotes the variance of an $\mathbb{R}$-valued random variable $Z\in{{\cal L}_{2}(\mathrm{P})}$. This motivates the following definition.
@@
for all $x_{1}^{\prime},x_{2}^{\prime}\in E^{\prime}$. The following lemma uses this observation to show that abstract covariance operators are determined by the behavior of $x^{\prime}\mapsto\langle x^{\prime},Tx^{\prime}\rangle_{E^{\prime},E}$, that is, on the diagonal.
@@
where the second identity can be found in e.g. [52, Theorem 1.1.2]. Moreover, a straightforward calculation shows that $\Lambda^{\dagger}$ is the diagonal $n\times n$-matrix having the entries $\lambda_{1}^{-1},\dots,\lambda_{n}^{-1}$ with $00^{-1}:=$ 0. Consequently, $A^{\dagger}$ is also symmetric and positive semi-definite. Moreover, this matrix describes the operator $T^{\dagger}:\mathbb{R}^{n}\to(\mathbb{R}^{n})^{\prime}$, and therefore $T^{\dagger}$ is symmetric and non-negative.
@@
Lemma B.2 shows that this kernel is determined by its behaviour on the diagonal. Moreover, the Cauchy-Schwarz inequality for kernels, see e.g. [42, (4.14)] yields
@@
Consequently, for every $x_{2}^{\prime}\in E^{\prime}$ with $\langle x_{2}^{\prime},Tx_{2}^{\prime}\rangle_{E^{\prime},E}=0$ we find $Tx_{2}^{\prime}=0$ by a simple application of Hahn-Banach’s theorem.
@@
To prove the latter, we fix some $x^{\prime}_{0}\in B_{E^{\prime}}$ as well as a sequence $(x_{n}^{\prime})\subset B_{E^{\prime}}$ such that $x_{n}^{\prime}\to x^{\prime}_{0}$ with respect to $w^{*}$. Moreover, we write $Z:=X-\mathbb{E}X$. By (B.4) and (B.1) we then find
@@
Let $E$ and $F$ be separable Banach spaces, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, and $X\in{{\cal L}_{2}(\mathrm{P},E)}$ and $Y\in{{\cal L}_{2}(\mathrm{P},F)}$ be centered. Moreover, let $S:F\to F^{\prime}$ be a bounded, linear, symmetric, and non-negative operator. Then the operator $K:E^{\prime}\to E$ defined by
@@
$\langle x_{2}^{\prime},Kx_{1}^{\prime}\rangle_{E^{\prime},E}$
$=\langle x_{2}^{\prime},\operatorname{cov}(X,Y)y_{1}^{\prime}\rangle_{E^{\prime},E}$
@@
$=\int_{\Omega}\int_{\Omega}\langle x_{1}^{\prime},X(\omega_{1})\rangle_{E^{\prime},E}\cdot\langle S\circ Y(\omega_{1}),Y(\omega_{2})\rangle_{F^{\prime},F}\cdot\langle x_{2}^{\prime},X(\omega_{2})\rangle_{E^{\prime},E}\,\mathrm{d}\mathrm{P}(\omega_{1})\,\mathrm{d}\mathrm{P}(\omega_{2})\,.$ (B.11)
@@
for all $\omega_{1},\omega_{2}\in\Omega$. Now, by assumption and the Cauchy-Schwarz inequality we know $\|X\|_{E}\cdot\|Y\|_{F}\in{{\cal L}_{1}(\mathrm{P})}$. This gives
@@
$\leq\|x_{1}^{\prime}\|\cdot\|x_{2}^{\prime}\|\cdot\|S\|\cdot\int_{\Omega}\|X(\omega_{1})\|_{E}\cdot\|Y(\omega_{1})\|_{F}\,\mathrm{d}\mathrm{P}(\omega_{1})\cdot\int_{\Omega}\|X(\omega_{2})\|_{E}\cdot\|Y(\omega_{2})\|_{F}\,\mathrm{d}\mathrm{P}(\omega_{2})$
@@
and therefore Fubini’s theorem applied to (B.11) yields the formula for $\langle x_{2}^{\prime},Kx_{1}^{\prime}\rangle_{E^{\prime},E}$.
@@
$(x^{\prime}\otimes y)(x):=\langle x^{\prime},x\rangle_{E^{\prime},E}\cdot y\,,\qquad\qquad x\in E\,.$ (B.14)
@@
for all $x^{\prime},x^{\prime}_{1},x^{\prime}_{2}\in E^{\prime}$, $y,y_{1},y_{2}\in F$, and $\alpha,\beta\in\mathbb{R}$, where, as usual, we assumed that $\otimes$ has a higher priority than $+$. We define the algebraic tensor product space by
@@
It is well-known that the set $\mathfrak{N}(E,F)$ of nuclear operators $T:E\to F$ equipped with $\|\,\cdot\,\|_{\mathrm{nuc}}$ is a Banach space and we have $\|T\|\leq\|T\|_{\mathrm{nuc}}$ for all $T\in\mathfrak{N}(E,F)$, see e.g. [54, Satz VI.5.3] or [14, Theorem 5.25]. Moreover, every finite rank operator is nuclear, and in particular all elementary tensors are nuclear. In addition, for $x^{\prime}\in E^{\prime}$ and $y\in F$ we have
@@
Under suitable assumptions, it turns out that the Bochner integral of tensor space valued functions can be used to compute cross covariances. To establish this result we need the following elementary lemma, which investigates such functions in the case of $X$ or $Y$ being constant.
@@
Proof of Lemma B.6: We only prove the assertion for $X^{\prime}\otimes y$, since the proof for $x^{\prime}\otimes Y$ is analogous.
Let us first assume that $X^{\prime}$ is a measurable step function. We fix a representation $X=\sum_{i=1}^{m}\boldsymbol{1}_{A_{i}}x_{i}^{\prime}$, where $A_{i}\in{\cal A}$ and $x_{i}^{\prime}\in E^{\prime}$. For $\omega\in\Omega$ the bilinearity of elementary tensors then gives
@@
$\|(X_{n}^{\prime}\otimes y)(\omega)-(X^{\prime}\otimes y)(\omega)\|_{\mathrm{nuc}}\to 0\,.$ (B.20)
Furthermore, (A.2) and (B.19) give $\|(X^{\prime}_{n}\otimes y)(\omega)\|_{\mathrm{nuc}}\leq\|(X^{\prime}\otimes y)(\omega)\|_{\mathrm{nuc}}$, and we already know that each $X^{\prime}_{n}\otimes y:\Omega\to\mathfrak{N}(E,F)$ is a measurable step function. In summary, $(X_{n}^{\prime}\otimes y)$ is a sequence approximating $X^{\prime}\otimes y$ in $\mathfrak{N}(E,F)$ in the sense of (A.1) and (A.2). Consequently, $X^{\prime}\otimes y:\Omega\to\mathfrak{N}(E,F)$ is strongly measurable as discussed around (A.1) and (A.2). Moreover, with the help of (B.18) we obtain
@@
Let us now turn to the announced integral representation of cross covariance operators. To this end, we fix two Banach spaces $E$ and $F$, as well as $x\in E$ and $y\in F$. For $x^{\prime\prime}:=\iota_{E}x\in E^{\prime\prime}$ we then write
@@
$\|x\otimes y\|$ $=\|x\otimes y\|_{\mathrm{nuc}}=\|x\|_{E}\cdot\|y\|_{F}\,,$ (B.21)
$\|x_{1}\otimes y_{1}\|_{\mathrm{nuc}}$ $\leq\|x_{2}\otimes y_{2}\|_{\mathrm{nuc}}$ (B.22)
for all $x,x_{1},x_{2}\in E$ and $y,y_{1},y_{2}\in F$ with $\|x_{1}\|_{E}\leq\|x_{2}\|_{E}$ and $\|y_{1}\|_{F}\leq\|y_{2}\|_{F}$. Moreover, Lemma B.5 also holds true, since $x_{n}\to x$ in $E$ implies $\iota_{E}x_{n}\to\iota_{E}x$ in $E^{\prime\prime}$.
@@
$\int_{\Omega}X\otimes y\,\mathrm{d}\mathrm{P}=\int_{\Omega}X^{\prime\prime}\otimes y\,\mathrm{d}\mathrm{P}=\biggr{(}\int_{\Omega}X^{\prime\prime}\,\mathrm{d}\mathrm{P}\biggl{)}\otimes y$ $=\biggr{(}\int_{\Omega}\iota_{E}\circ X\,\mathrm{d}\mathrm{P}\biggl{)}\otimes y$
$=\biggl{(}\iota_{E}\biggr{(}\int_{\Omega}X\,\mathrm{d}\mathrm{P}\biggl{)}\biggr{)}\otimes y$
$=\biggl{(}\int_{\Omega}X\,\mathrm{d}\mathrm{P}\biggr{)}\otimes y$ (B.23)
by an application of (A.10) and our nomenclature for tensors of the form $x\otimes y$. In particular, the Bochner integral of $X$ in (B.23) is an $E$-valued Bochner integral and not just an $E^{\prime\prime}$-valued integral as a naïve application of Lemma B.6 would suggest.
@@
where in the second to last step we used Lemma B.6 in the form of (B.23) twice.
@@
Furthermore, (A.2) and (B.22) gives $\|(X_{n}\otimes Y_{n})(\omega)\|_{\mathrm{nuc}}\leq\|(X\otimes Y)(\omega)\|_{\mathrm{nuc}}$ and we already know that each $X_{n}\otimes Y_{n}:\Omega\to\mathfrak{N}(E^{\prime},F)$ is a measurable step function. In summary, $(X_{n}^{\prime}\otimes Y_{n})$ is therefore a sequence approximating $X^{\prime}\otimes Y$ in $\mathfrak{N}(E,F)$ in the sense of (A.1) and (A.2). Consequently, $X\otimes Y$ is strongly measurable as discussed around (A.1) and (A.2). Moreover, with the help of (B.21) we obtain
@@
To establish (B.24), we first note that (A.1) and (A.2) for $X$ and $(X_{n})$ give $\mathbb{E}X_{n}\to\mathbb{E}X=0$ by the definition of Bochner integrals and our assumption that $X$ is centered. Analogously, we find $\mathbb{E}Y_{n}\to 0$. Combining both gives $\|\mathbb{E}X_{n}\otimes\mathbb{E}Y_{n}\|_{\mathrm{nuc}}\to 0$ by Lemma B.5. For $x^{\prime}\in E^{\prime}$ we thus find
@@
where in the first step we used the definition of Bochner integrals, the second step used that the limit in the second expression is with respect to the nuclear norm and thus also with respect to the operator norm, and in the last step we used the already established identity (B.24) for step functions. Moreover, we also have
@@
where the second step is based upon the dominated convergence property discussed around (A.7) to (A.9) with the $\mathrm{P}$-integrable $g:\Omega\to[0,\infty]$ defined by $g(\omega):=\|x^{\prime}\|_{E^{\prime}}\cdot\|X(\omega)\|_{E}\cdot\|Y(\omega)\|_{F}$ and the convergence
@@
for all $\omega\in\Omega$. Combining (B.25) and (B.26) yields (B.24). ∎
@@ C Appendix: Characteristic Functions
It is well known, see e.g. [29, Theorem 15.21], that every characteristic function $\varphi_{X}$ is uniformly continuous and we have $|\varphi_{X}(t)|\leq 1$ for all $t\in\mathbb{R}$ with $\varphi_{X}(0)=1$. Moreover, the characteristic function can be used to identify measures. The following well-known theorem, which can be found in e.g. [12, Proposition 10.3.11] or [29, Theorem 15.8] formalizes this.
@@
Proof of Lemma C.2: The second identity can be found in e.g. [25, Theorem 13.3], and the first identity is a special case of the second since $\langle a,X\rangle=a^{\mathsf{T}}X$ and $as=sa$, where $as$ denotes the matrix product and $a$ are $s$ are viewed as $d\times 1$ and $1\times 1$-matrices, respectively. ∎
@@
iii) $\Rightarrow$ i). See, for example, [16, Theorem 9.8.2] or [29, Theorem 15.23].
ii) $\Leftrightarrow$ iii). We first note that $\langle a,X_{n}\rangle\to\langle a,X\rangle$ in distribution is equivalent to $\varphi_{\langle a,X_{n}\rangle}(s)\to\varphi_{\langle a,X\rangle}(s)$ for all $s\in\mathbb{R}$ by the already established equivalence of i) and iii). Now the assertion quickly follows by applying the first formula of Lemma C.2. ∎
@@
Obviously, we have $a_{0}=|\varphi_{X}(t)|=1$. Since $\varphi_{X}$ is uniformly continuous there thus exists a $t_{0}>0$ such that $a_{t}>0$ for all $t\in[-t_{0},t_{0}]$. Consequently, we have
@@
see e.g. [50, Chapter IV.2.1]. Like in the finite dimensional case, the characteristic function can be used to identify probability measures. Namely, if $\mathrm{P}$ and $\mathrm{Q}$ are probability measures with $\varphi_{\mathrm{P}}=\varphi_{\mathrm{Q}}$, then we have $\mathrm{P}=\mathrm{Q}$, see e.g. [50, Theorem 2.2 in Chapter IV.2.1].
Similar to the finite dimensional case, a Gaussian measure $\mathrm{P}$ on a separable Banach space $E$ has a characteristic function that is solely described by its mean and covariance. The following result, which can be found in e.g. [50, Proposition 2.8 in Chapter IV.2.4], provides the details.
@@
Proof of Theorem C.6: We will apply [50, Theorem 3.1 in Chapter IV.3.1] for $\Gamma:=E^{\prime}$. Here we first note that Hahn-Banach’s theorem ensures that $E^{\prime}$ is a separating set consisting of continuous functions $E\to\mathbb{R}$. Moreover, every probability measure on $\mathfrak{B}(E)$ is a Radon probability measure, see the references around (46). Finally, [8, Theorem 3.8.4] in combination with i) ensures that $\{\mathrm{P}_{n}:n\geq 1\}$ is relatively weakly compact since the underlying space $E$ is a metric space, and thus a completely regular topological space, where the latter notion can be found in e.g. [8, Appendix A.1 on page 363]. Now the assertion follows from [50, Theorem 3.1 in Chapter IV.3.1]. ∎